OpenStudy (anonymous):

√10b+6=6 Help please showing work? still confused. ugh

OpenStudy (anonymous):

 confirm the problem

OpenStudy (anonymous):

\[\sqrt{10b+6}=6\]

OpenStudy (campbell_st):

square both sides of the equation and you get \[10b + 6 = 36\] now solve for b